### v1

# Generating patient cohorts from electronic health records using two-step retrieval-augmented text-to-SQL generation

Abstract
Clinical cohort definition is crucial for patient recruitment and observational studies, yet translating inclusion/exclusion criteria into SQL queries remains challenging and manual. We present an automated system utilizing large language models that combines criteria parsing, two-level retrieval augmented generation with specialized knowledge bases, medical concept standardization, and SQL generation to retrieve patient cohorts with patient funnels. The system achieves 0.75 F1-score in cohort identification on EHR data, effectively capturing complex temporal and logical relationships. These results demonstrate the feasibility of automated cohort generation for epidemiological research.
Generating patient cohorts from electronic health records using two-step retrieval-augmented text-to-SQL generation
Angelo Ziletti* Bayer AG angelo.ziletti@bayer.com                        Leonardo D’Ambrosi Bayer AG
1 Introduction
Electronic Health Records (EHR) have become essential resources for clinical research, serving as a foundational component across multiple healthcare applications. In clinical trials, accurate cohort identification directly impacts patient recruitment success, study validity, timeline adherence, and cost management Sherman et al. (2016). For observational studies and epidemiological research, well-defined cohorts enable researchers to investigate disease patterns, treatment effectiveness, and health outcomes across specific patient populations, informing critical healthcare decisions from drug development to regulatory submissions Casey et al. (2016); U.S. FDA (2024)
Problem Statement. Converting clinical inclusion/exclusion criteria into accurate database queries requires addressing multiple technical challenges simultaneously: (1) mapping natural language criteria to precise computational logic while leveraging medical domain knowledge, (2) preserving complex temporal and logical relationships across multiple criteria, (3) mapping diverse medical concepts to standardized codes, and (4) generating queries that conform to established data models.
Contributions. This paper presents an approach for automatically generating patient cohorts from EHRs. Our main contributions are:
  • A manually curated dataset of inclusion/exclusion criteria paired with SQL queries that conform to the Observational Medical Outcomes Partnership Common Data Model (OMOP-CDM) OMOP-CDM (2023).
  • A two-level retrieval-augmented generation (RAG) framework that substantially outperforms simple prompting across all tested models, demonstrating robust performance in translating complex clinical criteria into executable SQL queries.
  • Open-source release of dataset, source code, and prompt configurations [^1] Upon acceptance of the paper.
The system has been deployed and is currently being evaluated at Bayer by epidemiologists and data analysts.
Figure 1: (a-g) From inclusion/exclusion criteria in natural language to patient cohorts using electronic health record databases: end-to-end workflow. (h) Query decomposition and patient funnel generation through LLM-based processing.
2 Knowledge base generation
We create two complementary knowledge bases (KBs) for our retrieval-augmented approach: EpiAskKB and EpiCohoKB.
EpiAskKB consists of 111 question-SQL pairs that capture typical analytical questions in observational studies, extending and refining the dataset from Ziletti and DAmbrosi (2024) to better reflect common epidemiological research questions.
EpiCohoKB contains inclusion/exclusion criteria for observational studies, which we create through structured language model prompting. The generation integrates three key elements: (1) a prompt specifying OMOP CDM vocabulary usage and structural constraints for cohort criteria, (2) standard guidelines for study criteria (including temporal relationships, assessment windows, and observable time definitions) Wang et al. (2021), and (3) real observational study protocols from the European Medicines Agency (EMA) catalog EMA (2024) (89 samples) as exemplars for in-context learning. This ensures both structural rigor and clinical authenticity. EpiCohoKB covers diverse therapeutic areas and study types, including drug safety analyses, comparative effectiveness research, and healthcare utilization studies, creating a comprehensive benchmark for clinical criteria translation. We generate an initial set of SQL queries through large language model (LLM) prompting, manually review and validate them for Snowflake SQL compatibility, resulting in 104 samples.
3 Methods
Our approach translates clinical cohort definitions into executable SQL queries through a multi-stage pipeline (Fig. 1). First, we use an LLM to parse input criteria into a semi-structured format, explicitly identifying inclusion criteria, exclusion criteria, and index date definition. We then compile an instruction prompt incorporating domain knowledge about OMOP CDM and SQL query construction, forming our baseline zero-shot (ZS) approach.
We implement three retrieval augmentation strategies: (1) Criteria-level RAG (RAG+A) retrieves top-5 similar question-SQL pairs from EpiAskKB per criterion; (2) Cohort-level RAG (RAG+C) retrieves top-5 similar cohort definitions from EpiCohoKB; (3) Combined RAG (RAG+A+C) performs both criteria-level and cohort-level retrieval. This modular design enables query pattern reuse at different granularities - individual criteria can be composed into complex cohort definitions while preserving temporal relationships. The augmented prompt is passed to the LLM to generate SQL with placeholder entities (e.g., condition@hypertension). This approach disentangles the query logic generation from the subsequent medical coding step where placeholders are mapped to specific concept IDs Ziletti and DAmbrosi (2024). The system implements a self-healing mechanism with up to three iterations Pourreza and Rafiei (2023). The output is a cohort table with patient IDs and index dates.
To provide interpretability and enable criteria refinement, we also implement a patient funnel calculation pipeline (Fig. 1h). The LLM generates separate SQL queries per criterion, which are executed independently and combined through set operations (intersections for inclusions, differences for exclusions), tracking cohort attrition at each step.
4 Evaluation
4.1 Experimental setup
[…]
4.2 Experimental results
Results are shown in Table 1, and outlined below.
Model name Settings General metrics Patient-level metrics Date-level metrics
Valid SQL Retrieved F1 Prec. Recall Size sim. Date overlap Within 30d
Claude 3.5 Sonnet ZS 100.0 78.1 57.8 61.0 63.1 62.8 75.9 74.9
Claude 3.5 Sonnet RAG+A 99.0 82.6 58.4 59.5 68.6 62.6 80.8 77.5
Claude 3.5 Sonnet RAG+C 99.0 91.4 69.5 72.0 74.7 74.5 88.2 87.4
Claude 3.5 Sonnet RAG+A+C 99.0 93.3 72.8 75.6 79.0 77.6 91.7 90.5
Gemini 2.0 Flash ZS 96.2 55.2 42.4 43.3 44.5 46.6 51.7 50.9
Gemini 2.0 Flash RAG+A 99.0 62.9 41.6 42.1 48.6 46.0 58.7 58.4
Gemini 2.0 Flash RAG+C 99.0 89.5 68.0 68.9 75.2 73.5 85.4 84.6
Gemini 2.0 Flash RAG+A+C 99.0 91.3 75.4 76.6 79.7 79.8 89.1 89.1
GPT-4o ZS 84.8 61.0 35.4 38.1 39.7 42.1 52.9 52.5
GPT-4o RAG+A 86.7 50.5 31.7 32.8 34.7 37.4 44.9 44.6
GPT-4o RAG+C 97.1 91.4 72.1 73.1 76.8 77.7 87.0 87.9
GPT-4o RAG+A+C 98.1 84.8 65.7 67.7 71.1 70.5 80.9 82.3
o1 ZS 9.5 5.7 4.3 5.2 4.2 4.5 5.4 5.0
o1 RAG+A+C 78.1 73.3 60.0 62.0 62.2 63.0 71.5 70.0
DeepSeek R1 ZS 35.2 28.6 22.2 22.6 23.1 23.6 25.9 24.8
DeepSeek R1 RAG+A+C 36.2 31.4 23.6 25.0 25.4 25.4 29.6 30.2
LLAMA 3.1 70B ZS 37.1 17.1 8.5 9.6 9.0 11.0 16.2 14.5
LLAMA 3.1 70B RAG+A+C 79.1 69.5 47.0 51.2 52.0 52.7 66.6 65.3
Table 1: Performance evaluation of text-to-SQL generation for patient cohort identification. Valid SQL indicates syntactically correct queries, Retrieved indicates queries that successfully retrieved patient data. Patient-level metrics evaluate cohort membership accuracy, while date-level metrics assess the temporal alignment of cohort index dates. Higher values are better, all values in percentage. Bold indicates best results, underlined shows second best.
Generated queries have high syntactic validity. Most models show high SQL compilation rates with RAG augmentation, while also improving data retrieval success. The retrieved data rate increases substantially with RAG.
RAG strongly improves zero-shot performance. The combined RAG strategy (RAG+A+C) substantially outperforms ZS across all models. RAG+A+C significantly improves patient cohort identification across all metrics. Precision increases (Claude: $\uparrow$14.6pp, Gemini $\uparrow$33.3pp, GPT-4o $\uparrow$29.6pp), while maintaining high recall (>70%). Similar patterns emerge in temporal alignment. The balanced improvement in precision and recall suggests that RAG helps models better understand both inclusion and exclusion criteria.
Complementary RAG strategies enhance performance. Cohort-level examples (RAG+C) are particularly effective. When compared to criteria-level RAG (RAG+A), cohort-level examples lead to substantially higher F1 scores. The combination of both strategies (RAG+A+C) further improves performance for most models, suggesting complementary benefits from both example types. Interestingly, GPT-4o shows slightly decreased performance with RAG+A+C compared to RAG+C ($\downarrow$6.4pp), indicating that longer input contexts may impact its generation capabilities.
Claude and Gemini emerge as top performers. Claude and Gemini achieve comparable performance when augmented with RAG+A+C. Claude exhibits the strongest ZS performance, while Gemini shows excellent in-context learning capabilities ($\uparrow$33.3pp from ZS to RAG+A+C). Reasoning-focused models (o1, DeepSeek R1) struggle with SQL formatting despite their analytical capabilities.
To evaluate the patient funnel creation, for the best performing model (Claude RAG+A+C), we measure a 90% normalized cohort size similarity compared to the single-query approach, validating the funnel decomposition strategy.
5 Related Work
Cohort Generation from EHR Data. Rule-based systems achieve high precision but require substantial expert curation Banda et al. (2017); Hripcsak et al. (2019); OHDSI Collaborative (2021); Privitera and Hartenstein (2024). While early work combined rules and machine learning Yuan et al. (2019), recent approaches leverage LLMs: Park et al. (2024) for OMOP-CDM clinical trial queries, Melnichenko (2023) for structured criteria extraction, Dobbins et al. (2023) for cross-schema templated queries, Yan et al. (2024) for phenotyping, and Wong et al. (2023) for eligibility criteria structuring. Ziletti and DAmbrosi (2024) introduced RAG-enhanced epidemiological text-to-SQL. Several EHR text-to-SQL datasets exist Wang et al. (2020); Raghavan et al. (2021); Lee et al. (2022); Tarbell et al. (2023); Ziletti and DAmbrosi (2024), though none maps observational study criteria to OMOP-SQL.
Text-to-SQL with In-Context Learning. Recent work shows LLMs excel at text-to-SQL tasks when enhanced with in-context learning Rajkumar et al. (2022), particularly through optimized example selection Nan et al. (2023); Gao et al. (2023), query decomposition Pourreza and Rafiei (2023), diverse demonstrations Chang and Fosler-Lussier (2023a); Chang and Fosler-Lussier (2023b), ensemble approaches Gao et al. (2025), and chain-of-thought prompting Zhang et al. (2023a).
6 Conclusions
This work shows that LLMs with RAG can effectively identify patient cohorts from EHR data. The pipeline combines medical concept standardization, criteria parsing, and two-level retrieval to handle complex temporal relationships, while the patient funnel provides interpretable intermediate metrics. Future applications to clinical trial recruitment could help bridge the gap between observational and interventional research.
7 Limitations
Our work has several limitations. The modest size of our dataset may not capture the full complexity and variety of real-world observational studies. The current system lacks formal uncertainty estimates for its predictions, making it difficult to identify cases where human review is particularly important. While the patient funnel provides some transparency, the model remains largely a black box for users unfamiliar with SQL, potentially limiting adoption among clinical researchers. The current methodology relies on OMOP-CDM, and although it could be in principle adapted to other data models through in-context learning, this was not tested.